Two resistors x and y are connected in series with a source voltage of 12v. X has a resistance of 4 ohm and the voltage drop across y is 8 v. Find the resiance of y and the current.

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
Voltage drop across y = 8 volts
Hence voltage drop across x = 12- 8 = 4 volts
Now current through x = 4/4 = 1 amp.
But x and y are in series hence current through y = 1 amp
Hence resistance across y = 8/1 = 8 ohm